# Identification and estimation of vaccine effectiveness in the test-negative design under equi-confounding

### v2

@@ 5.2 Identification in a cohort study
Given Assumptions A1 - A4, $\Psi$ is identifiable under cohort sampling, i.e. $O_{cohort}=\{(X_{i},V_{i},Y_{i}):i=1,\ldots,N\},$ by the expressions
@@
The first expression is a generalization of the result for $\Psi_{om}(X)$ discussed earlier, where the de-biasing term is now standardized over the distribution of $X$ among the vaccinated. The second expression offers an alternative representation of the de-biasing term based on the invariance property of the odds ratio. By combining this property with Assumptions A1 - A4, we can show that expression 2 is equivalent to
@@
Thus, the odds of vaccination among the test-negative controls may be viewed as another proxy for the de-biasing term. Following standard terminology in causal inference [robins_estimating_1992], we refer to expression A8 as the outcome-modeling estimand and expression A9 as the inverse-probability weighting estimand.
@@
The quantity $\Pr[V=1|Y^{0}=2,X]$ is sometimes referred to as the extended propensity score (EPS) function, as it extends the standard propensity score to allow for unmeasured confounding by conditioning on the treatment-free potential outcome — in this case $Y^{0}=2$. As shown in previous work [tchetgen_single_2023], in the presence of unmeasured confounding, knowledge of the EPS is sufficient to identify treatment effects among the treated. Similarly knowledge of $\Pr[V=1|Y^{1}=2,X]$ would permit identification of treatment effects among the untreated, and knowledge of both would permit the identification of the overall effect [tchetgen_single_2023].
@@ 5.3 Identification under outcome-dependent sampling in a TND
As with $\Psi(X)$ above, expressions A8 and A9 require knowledge of the absolute probability of receiving a test. However, in a TND we observe only individuals who pass the symptom screen and are tested. Despite this, the next proposition shows that $\Psi$ remains identifiable under the outcome-dependent sampling scheme of the TND.
@@
Under test-negative sampling, i.e. $O_{TND}=\{(X_{i},V_{i},S_{i}=1,Y_{i}):i=1,\ldots,n\},$ with selection $S=\mathbbm{1}(Y\neq 0)$, $\Psi_{om}$ and $\Psi_{ipw}$ are equivalent to
@@
Expressions 7 and 8 condition on selection in the TND and are thus identifiable based on the observed data. As before, we refer to expression 7 as the outcome-modeling estimand of the TND and expression 8 as the inverse-probability weighting estimand of the TND.
@@
Comparing expressions A8 and 7, we note that the “sampling scheme” of the TND can be viewed as outcome-dependent or case-control sampling of an outcome of interest ($Y=2$) and a corresponding negative outcome control ($Y=1$).
@@
Expressions 7 and 8 resemble those proposed assuming unconfoundedness [jiang_tnddr_2023]. In that work, the authors introduce a class of marginal estimands in a TND that can be written as
@@
where $\omega_{v}(x)$ are “de-biasing weights” used to adjust for bias due to outcome-dependent sampling. They describe weights based on both outcome-modeling and inverse probability weighting. We note that the denominators of expressions 7 and 8 both have a similar form, using alternative weights
@@ 6 Estimation
We now describe estimation strategies based on the identification results above in a TND. In the main text, we focus on the results from Proposition 2, but include in Appendix C estimators based on $\Psi_{om}^{*}$ and $\Psi_{ipw}^{*}$ Proposition 1 that, in theory, could be applied to the full cohort in the event that testing probabilities are known. For convenience, we define an indicator for test positive case in the TND sample $Y^{*}=\mathbbm{1}(Y=2)$.
The traditional approach to estimating vaccine effectiveness in a TND involves logistic regression of $Y$ conditional on $X$ and $V$, such as:
@@
where effectiveness is defined as one minus the odds ratio comparing vaccinated and unvaccinated, i.e. $VE=1-\exp(\gamma_{1})$. As we have shown, $\exp(\gamma_{1})$ is equal to $\Psi^{*}_{om}(X)$, and by extension, the conditional causal risk ratio among the vaccinated, $\Psi(X)$, under identifiability Assumptions A1 - A4 and correct model specification.
Alternatively, expressions 7 and 8 suggest two plug-in estimators valid under effect modification. Expression 7 suggests the plug-in estimator:
@@
Expression 8, on the other hand, suggests the plug-in estimator:
@@
The first estimator, $\widehat{\Psi}_{om}^{*}$, requires correct specification of the outcome model, while the second estimator, $\widehat{\Psi}_{ipw}^{*}$, requires correct specification of the extended propensity score model. The preferred estimator may depend on the context. For instance, if more is known about the assignment mechanism for vaccination, the weighting estimator may be preferred [robins_estimating_1992, braitman_rare_2002]. Conversely, if the process leading to infection and testing is better understood, the outcome estimator may be more appropriate. In practice, however, both models may be difficult to specify correctly. To address this, we construct a doubly-robust estimator for $\Psi$ that is consistent and asymptotically normal if the model for either $\mu^{*}_{0}(X)$ or $\pi^{*}_{0}(X)$ or both are correctly specified. Full details of the doubly robust approach are included in Appendix E, but are based on noting that an equivalent expression of $\Psi_{om}$ and $\Psi_{ipw}$ is
@@ 8 Simulation
To illustrate a data generation process where our assumptions hold and to evaluate the empirical performance of the proposed estimators when they do not, we conduct a simulation study mimicking a TND with binary vaccination, testing, and infection. The data generation process is as follows:
@@
This implies the conditional independence $Y^{v}\operatorname{\perp\!\!\!\perp}V\mid X,U$ holds for $v=0,1$, meaning that conditioning on $U$ and $X$ is sufficient to control confounding. When there is no effect modification by $U$ or $X$, the causal risk ratio for medically-attended illness is
$\Psi$ $=\exp(\beta_{2V}+\tau_{2V}).$
@@
We generate a target population of $N=15,000$ resulting in TND samples of test-positive cases and test-negative controls of between $1000$ and $2000$. We consider nine scenarios:
@@
Parameter values for all scenarios are shown in Table A1. For each scenario, we generate 1000 replicates and estimate $\Psi$ using the proposed estimators $\widehat{\Psi}_{om}^{*}$, $\widehat{\Psi}_{ipw}^{*}$, and $\widehat{\Psi}_{dr}^{*}$ as well as the conventional logistic regression estimator and calculate the bias and coverage of 95% confidence intervals based on the true value. For comparison, we also estimate $\Psi$ assuming one had access to the full sample, as in cohort study. We consider both the typical case that the covariate $U$ is unmeasured, and the hypothetical case where $U$ is available, allowing for conditioning on a sufficient adjustment set. Finally, we include estimates from and $\widehat{\Psi}_{om}$ based on $\Psi_{om}$ from Proposition 1 that uses the full sample instead of restricting to those tested. Finally, in scenario 9, we demonstrate the double robustness property of $\widehat{\Psi}_{dr}^{*}$ by adjusting the data generation process for $V$ and $I^{v}$ to include squared terms in $X$. Additional details on the setup and precise estimation methods are provided in Appendix I.
The results are reported in Figures 2 and 3 with further details in Tables A3 and A4. When there is no unmeasured confounding, all estimators are unbiased. In all other scenarios the typical cohort estimator without $U$ is biased (scenario 1). When there is unmeasured confounding but our assumptions hold, the TND estimator is unbiased (scenario 2). When there is an effect of vaccination on the test-negative illness (scenario 3), the TND estimator is biased towards the null (when the effect is also protective); however, it does identify the ratio of vaccine effects against $I=2$ versus $I=1$. When equi-confounding or equi-selection are violated the TND estimator is biased (scenarios 4 and 5). When the test-negative and test-positive illnesses are not mutually exclusive, the TND estimator is biased for the causal risk ratio but unbiased for the causal odds ratio. When there is a direct effect of vaccination on testing behavior, the TND estimator remains unbiased if the effect is equivalent for test-negative and test-positive illnesses (scenario 6), but is biased when they are unequal (scenario 7). Finally, the proposed TND estimators, $\widehat{\Psi}_{om}^{*}$, $\widehat{\Psi}_{ipw}^{*}$, and $\widehat{\Psi}_{dr}^{*}$, remain unbiased in the presence of effect modification while the conventional logistic regression approach is biased for the marginal estimand (scenario 9). As expected, when the outcome model is misspecified $\widehat{\Psi}_{ipw}^{*}$ and $\widehat{\Psi}_{dr}^{*}$ remain unbiased while $\widehat{\Psi}_{om}^{*}$ is biased. Likewise, when the extended propensity score model is misspecified $\widehat{\Psi}_{om}^{*}$ and $\widehat{\Psi}_{dr}^{*}$ remain unbiased while $\widehat{\Psi}_{ipw}^{*}$ is biased. When both models are misspecified all estimators are biased.
@@ 9 Discussion
In this article, we described alternative conditions under which the TND produces an unbiased estimate of vaccine effectiveness in the presence of unmeasured confounding. We demonstrated that the TND accurately recovers the vaccine effectiveness among the vaccinated, provided that the vaccine does not affect the test-negative illness and that unmeasured variables influence symptomatic illness and testing equivalently for both test-positive and test-negative infections. While our results apply to commonly-used logistic regression estimator, we also derived novel estimators based on outcome modeling, inverse probability weighting, and the efficient influence function.
@@
The plausibility of the equi-confounding assumption depends on understanding the source of the test-negative illness, which requires additional sampling and testing. In previous TNDs for influenza the most common test-negative illnesses were rhinovirus, respiratory syncytial virus (RSV), human metapneumovirus (hMPV), parainfluenza virus (PIV), bocavirus (BoV), coronavirus (CoV), adenovirus (AdV), enterovirus (EV) and others [chua_use_2020-1]. Ideally, the test-negative illness should have a similar exposure mechanism and clinical presentation as the target illness but remain unaffected by the vaccine. For example, in pneumococcal disease, non-vaccine serotypes of Streptococcus pneumoniae serve as ideal test-negative controls [broome_pneumococcal_1980]. In Appendix H, we explore potential structural sources of equi-confounding —including health-seeking behavior and immortal time— as well as example violations, such as correlated vaccine-seeking for vaccines targeting test-negative pathogens [payne_impact_2023] and at home testing [qasmieh2024magnitude] and how these may be resolved. In practice, assessing the appropriateness of the equi-confounding assumption will require subject-matter expertise and should be evaluated on a case-by-case basis.
@@
Even when equi-confounding is not plausible, as our simulations suggest, the TND may still yield less biased estimates compared to conventional approaches based on a minimal covariate set, especially when the test-negative illness serves as a reasonable, albeit imperfect, proxy for residual confounding. Our approach also permits a straightforward sensitivity analysis that can be used to understand the implications of departures from equi-confounding. Alternatively, when equi-confounding is implausible, additional negative controls can be leveraged, as suggested by recent methods using multiple proxies for unmeasured confounding [li_double_2023].
@@
Our results also depend on the assumption that the vaccine does not affect other infections or influence test-seeking behavior (except through infection). The former may be violated if either vaccination provides short-term nonspecific protection through immune activation. Previous research has suggested that this is possible for influenza vaccines [cowling_increased_2012]. However, other results have shown the opposite [sundaram_influenza_2013]. Violations of this assumption would bias TND estimates, though in some cases, if the effect is in the same direction, the TND might still provide a lower bound on true vaccine effectiveness. If vaccination exerts equal effects on test-seeking behavior for test-positive and test-negative infections, as our simulations demonstrate, estimates of vaccine effectiveness against symptomatic infection remain unbiased. Conversely, when effects are unequal, the TND estimate could be biased.
@@
Figure 2: Bias of vaccine effectiveness (VE) estimators across different simulation scenarios. In each scenario, we compare: (i) ideal cohort estimator with $U$ measured, (ii) more typical cohort estimator with $U$ unmeasured, (iii) a difference-in-differences (DiD) estimator in the cohort based on equation A8, and (iv) the TND estimator $\widehat{\Psi}_{om}^{*}$. Results are based on 1000 Monte Carlo replications. Black dashed line is the true VE. In scenario 3, grey dashed line is the ratio of VEs for test-positive versus test-negative illness. In scenarios 6 and 7, grey dashed line VE against symptomatic illness.
@@ Appendix A A nonparametric odds ratio model
$\Pr(Y^{0}=2|V=1,X=x)=\dfrac{\Pr(Y^{0}=2|V=1,X=x)\exp\{\beta(2,x)\}}{\int f(Y^{0}=y|V=0,X=x)\exp\{\beta(y,x)\}\;dy}.$

### v3

@@ 5.2 Identification in a cohort study
Given Assumptions A1 - A4, $\Psi$ is identifiable under cohort sampling, i.e. $O_{cohort}=\{(X_{i},V_{i},Y_{i})\}_{i=1}^{N},$ by the expressions
@@
Expression 6 is a generalization of the result for $\Psi_{om}(X)$ discussed earlier, where the de-biasing term is now standardized over the distribution of $X$ among the vaccinated. Expression 7 offers an alternative representation of the de-biasing term based on Assumptions A1 - A4 and the invariance property of the odds ratio, as
@@
Following standard terminology [robins_estimating_1992], we refer to Expression 6 as the outcome-modeling estimand and Expression 7 as the inverse-probability weighting estimand.
@@
The quantity $\Pr[V=1|Y^{0}=2,X]$ is sometimes referred to as the extended propensity score (EPS) function, as it extends the standard propensity score to allow for unmeasured confounding by conditioning on the treatment-free potential outcome — in this case $Y^{0}=2$. As shown previously [tchetgen_single_2023], in the presence of unmeasured confounding, knowledge of the EPS is sufficient to identify treatment effects among the treated. Similarly knowledge of $\Pr[V=1|Y^{1}=2,X]$ would permit identification of treatment effects among the untreated, and knowledge of both would permit the identification of the overall effect [tchetgen_single_2023].
@@ 5.3 Identification under outcome-dependent sampling in a TND
Expressions 6 and 7 require absolute probabilities that cannot be estimated under the outcome-dependent sampling scheme of the TND. Nonetheless, the next proposition shows that $\Psi$ remains identifiable
@@
Under test-negative sampling, i.e. $O_{TND}=\{(X_{i},V_{i},S_{i}=1,Y_{i})\}_{i=1}^{n},$ with selection $S=\mathbbm{1}(Y\neq 0)$, $\Psi_{om}$ and $\Psi_{ipw}$ are equivalent to
@@
Expressions 8 and 9 only require conditional probabilities among the tested and thus are identifiable in a TND. As before, we refer to Expression 8 as the outcome-modeling estimand of the TND and Expression 9 as the inverse-probability weighting estimand of the TND.
@@
Comparing Expressions 6 and 8, we note that the “sampling scheme” of the TND can be viewed as outcome-dependent or case-control sampling of an outcome of interest ($Y=2$) and a corresponding negative outcome control ($Y=1$).
@@
Expressions 8 and 9 resemble those proposed assuming unconfoundedness [jiang_tnddr_2023]. In that work, the authors introduce a class of marginal estimands in a TND that can be written as
@@
where $\omega_{v}(x)$ are “de-biasing weights” used to adjust for bias due to outcome-dependent sampling. They describe weights based on both outcome-modeling and inverse probability weighting. We note that the denominators of Expressions 8 and 9 both have a similar form, using alternative weights
@@ 6 Estimation
We now describe estimation strategies for the causal risk ratio among the vaccinated in a TND, based on the identifying expressions in Proposition 2. However, we describe in Appendix C estimators based on the results in Proposition 1 in a cohort study. For convenience, we define an indicator for test positive case in the TND sample $Y^{*}=\mathbbm{1}(Y=2)$.
Traditionally in a TND, the analyst performs a logistic regression of $Y$ conditional on $X$ and $V$, such as:
@@
where $\gamma_{1}$ is the focal coefficient. As we have shown, $\exp(\gamma_{1})$ is equal to $\Psi^{*}_{om}(X)$, and by extension, the conditional causal risk ratio among the vaccinated, $\Psi(X)$, under identifiability Assumptions A1 - A4 and correct model specification.
Alternatively, Expressions 8 and 9 suggest two plug-in estimators valid under effect modification. Expression 8 suggests the plug-in estimator:
@@
Expression 9, on the other hand, suggests the plug-in estimator:
@@
The first estimator, $\widehat{\Psi}_{om}^{*}$, requires correct specification of the outcome model, while the second estimator, $\widehat{\Psi}_{ipw}^{*}$, requires correct specification of the extended propensity score model. In practice, however, both models may be difficult to specify correctly. To address this, we construct a doubly-robust estimator for $\Psi$ that is consistent and asymptotically normal if the model for either $\mu^{*}_{0}(X)$ or $\pi^{*}_{0}(X)$ or both are correctly specified. Full details of the doubly robust approach are included in Appendix E, but are based on noting that an equivalent expression of $\Psi_{om}$ and $\Psi_{ipw}$ is
@@ 8 Simulation
To illustrate a data generation process where our assumptions hold and to evaluate the empirical performance of the proposed estimators when they do not, we conduct a simulation study. The data generation process is as follows:
@@
This implies the conditional independence $Y^{v}\operatorname{\perp\!\!\!\perp}V\mid X,U$ holds for $v=0,1$, meaning that conditioning on $U$ and $X$ is sufficient to control confounding. When there is no effect modification by $U$ or $X$ ($\beta_{2VX}=0$) and no direct effect on testing for focal illness ($\tau_{2V}=0$), the causal risk ratio is
$\Psi$ $=\exp(\beta_{2V}).$
@@
Likewise, violations of Assumption A2 are controlled by $\beta_{1V}$ and $\tau_{1V}$. When $\beta_{1V}\neq 0$ there is a direct effect of vaccination on test-negative illness $I=1$ and when $\tau_{1V}\neq 0$ there is a direct effect of vaccination on testing among those with test-negative illness. When this occurs, we can at most identify the ratio of vaccine effects on the two illness outcomes, i.e. $\exp(\beta_{2V}+\tau_{2V})/\exp(\beta_{1V}+\tau_{1V})$, provided there is no effect modification by $X$ or $U$. As we show in Appendix F and illustrate further via simulation below, under the special case of equal effects on testing, i.e. $\tau_{2V}=\tau_{1V}$, certain effects on symptomatic illness may still be recovered.
We generate a target population of $N=15,000$ resulting in TND samples of test-positive cases and test-negative controls of between $1000$ and $2000$. We consider eight scenarios:
@@
Parameter values for all scenarios are shown in Table A1. For each scenario, we generate 1000 replicates and estimate $\Psi$ using the proposed estimators $\widehat{\Psi}_{om}^{*}$, $\widehat{\Psi}_{ipw}^{*}$, and $\widehat{\Psi}_{dr}^{*}$ as well as the conventional logistic regression estimator and calculate the bias and coverage of 95% confidence intervals based on the true value. For comparison, we also consider the following estimators based on data from the full cohort: a conventional outcome-modeling estimator adjusting for $X$ alone, another conditioning on $X$ and $U$ (covering the hypothetical case that all confounders are measured), and one based on $\Psi_{om}$ in Proposition 1. Finally, in scenario 8, we demonstrate the double robustness property of $\widehat{\Psi}_{dr}^{*}$ by adjusting the data generation process for $V$ and $I^{v}$ to include squared terms in $X$. Additional details on the setup and precise estimation methods are provided in Appendix I.
The results are reported in Figures 2 and 3 with further details in Tables A3 and A4. When there is no unmeasured confounding, all estimators are unbiased. In all other scenarios the typical cohort estimator without $U$ is biased (scenario 1). When there is unmeasured confounding but our assumptions hold, the TND estimator is unbiased (scenario 2). When there is an effect of vaccination on the test-negative illness (scenario 3), the TND estimator is biased towards the null (when the effect is also protective); however, it does identify the ratio of vaccine effects against $I=2$ versus $I=1$. When equi-confounding or equi-selection are violated the TND estimator is biased (scenarios 4 and 5). When the test-negative and test-positive illnesses are not mutually exclusive, the TND estimator is biased for the causal risk ratio but unbiased for the causal odds ratio. When there is a direct effect of vaccination on testing behavior, the TND estimator remains unbiased if the effect is equivalent for test-negative and test-positive illnesses (scenario 6), but is biased when they are unequal (scenario 7). Finally, the proposed TND estimators, $\widehat{\Psi}_{om}^{*}$, $\widehat{\Psi}_{ipw}^{*}$, and $\widehat{\Psi}_{dr}^{*}$, remain unbiased in the presence of effect modification while the conventional logistic regression approach is biased for the marginal estimand (scenario 8). As expected, when the outcome model is misspecified $\widehat{\Psi}_{ipw}^{*}$ and $\widehat{\Psi}_{dr}^{*}$ remain unbiased while $\widehat{\Psi}_{om}^{*}$ is biased. Likewise, when the extended propensity score model is misspecified $\widehat{\Psi}_{om}^{*}$ and $\widehat{\Psi}_{dr}^{*}$ remain unbiased while $\widehat{\Psi}_{ipw}^{*}$ is biased. When both models are misspecified all estimators are biased.
@@ 9 Discussion
We have shown that the TND can yield unbiased estimates of vaccine effectiveness among the vaccinated in the presence of unmeasured confounding provided certain conditions are met: namely, the vaccine does not affect the test-negative illness and the net effect of unmeasured confounders on symptomatic illness and testing is equal for both test-positive and test-negative infections. While our results apply to commonly-used logistic regression estimator, we also derived novel estimators based on outcome modeling and inverse probability weighting as well as a doubly-robust estimator that is consistent if one or both of the outcome model and propensity score model are correctly specified.
@@
The plausibility of the equi-confounding assumption may be increased by knowledge of the source of the test-negative illness. In previous TNDs for influenza that included additional testing to confirm the source of test-negative illness the most common causes were rhinovirus, respiratory syncytial virus (RSV), human metapneumovirus (hMPV), parainfluenza virus (PIV), bocavirus (BoV), coronavirus (CoV), adenovirus (AdV), enterovirus (EV) and others [chua_use_2020-1]. Ideally, the test-negative illness should have a similar exposure mechanism and clinical presentation as the target illness but remain unaffected by the vaccine. For example, in pneumococcal disease, non-vaccine serotypes of Streptococcus pneumoniae serve as ideal test-negative controls [broome_pneumococcal_1980]. In Appendix H, we explore potential structural sources of equi-confounding —including health-seeking behavior and immortal time— as well as example violations, such as correlated vaccine-seeking [payne_impact_2023] and at home testing [qasmieh2024magnitude] and how these may be resolved.
@@
Even when equi-confounding does not hold, as our simulations suggest, the TND may still reduce bias relative to other approaches based only on a minimal covariate set, when test-negative illness serves as a reasonable, if imperfect, proxy for residual confounding. We also propose a sensitivity analysis that allows for straightforward departures from equi-confounding via the exponential tilt model. Alternatively, additional negative controls can be leveraged in a proximal inference framework to identify effects under more general unmeasured confounding structures [li_double_2023].
@@
Our results also depend on the assumption that the vaccine does not affect other infections or influence test-seeking behavior (except through infection). The former may be violated if vaccination provides short-term nonspecific protection through immune activation. Previous research has suggested that this is possible for influenza vaccines [cowling_increased_2012]. However, other results have shown the opposite [sundaram_influenza_2013]. Violations of this assumption would bias TND estimates, though in some cases, if the effect is in the same direction, the TND might still provide a lower bound on true vaccine effectiveness. If vaccination exerts equal effects on test-seeking behavior for test-positive and test-negative infections, as our simulations demonstrate, estimates of vaccine effectiveness against symptomatic infection remain unbiased.
@@
Figure 2: Bias of vaccine effectiveness (VE) estimators across different simulation scenarios. In each scenario, we compare: (i) ideal cohort estimator with $U$ measured, (ii) more typical cohort estimator with $U$ unmeasured, (iii) a difference-in-differences (DiD) estimator in the cohort based on equation 6, and (iv) the TND estimator $\widehat{\Psi}_{om}^{*}$. Results are based on 1000 Monte Carlo replications. Black dashed line is the true VE. In scenario 3, grey dashed line is the ratio of VEs for test-positive versus test-negative illness. In scenarios 6 and 7, grey dashed line VE against symptomatic illness.
@@ Appendix A A nonparametric odds ratio model
$\Pr(Y^{0}=2|V=1,X=x)=\dfrac{\Pr(Y^{0}=2|V=0,X=x)\exp\{\beta(2,x)\}}{\int f(Y^{0}=y|V=0,X=x)\exp\{\beta(y,x)\}\;dy}.$